For a certain gas enclosed in a container of fixed volume, the pressure P (in newtons per square
meter) is directly proportional to the temperature T (in Kelvin). If the pressure is found to be 100
newtons per square meter at a temperature of 300 Kelvin, find the equation of variation. Find the
temperature when the pressure is 120 newtons per square meter.

Answer by Alan3354(69443)

For a certain gas enclosed in a container of fixed volume, the pressure P (in newtons per square meter) is directly proportional to the temperature T (in Kelvin). If the pressure is found to be 100 newtons per square meter at a temperature of 300 Kelvin, find the equation of variation. Find the temperature when the pressure is 120 newtons per square meter.
====================
P = k*T --- k is a constant
---
k = P/T = 1/3
===================
T = P/k = 3P
T = 360K